20:30: we point telescope

LAN Asti ToF
-160k 460 84.0
FLAO Z stage is at 50mm, we change the ToF to 84us and we close the loop. We change in closed loop the ToF to 85.8us (the correct value for LUCI focus) but the system crashes.

So we close the loop on FLAO, 10 modes, stage at 27mm, ToF 85.8us.

LAN Asti ToF
-110k 410 85.8

TN Notes
20141201_051941 G=2.0,0.5,0.4
We repoint the telescope on globular cluster, we set the FLAO on Bin4.

21:45: we implement the autogain: optimal gains are 10,0.6, 0.4 but trying to close the loop with these gains the system crashes.

TN Notes
20141201_061601 seeing=*1.2"*, G=5.0,0.4,0.3
22:15: we reoptimize the LGS spot size.

LAN Asti ToF
-180k 960 85.8

Vibration compensation coordinates system

We take several snapshots enabling the Vibration compensation system and rising the gain to make the system oscillate:

Vib TN
OFF 20141201_063816
LM1 phi -0.5 20141201_063919
LM1 phi 0.5 20141201_064125
LM1 theta 20141201_064532
LM2 phi 20141201_064633
LM2 theta 20141201_064721

GLAO operation

we go back to the GLAO loop operation, we close 100modes on ARGOS:

TN Notes
20141201_065717 G=5.0,0.6,0.4
20141201_065935 OL, seeing=*1.5"*
We run the autogain procedure. We run the TS on 10 modes but the system oscillates, we lower the gain on the first 36 modes to 0.5. We also increase the opening time to 1500ns to have ~4k ADU per subap.

TN Notes
20141201_070836  
20141201_071217 G=5.0,0.3,0.1
20141201_071255 stage 27mm, elevation 66°
20141201_071358 stage 30mm
20141201_071555 stage 35mm
20141201_071638 stage 40mm
20141201_072354 G=5.0,0.6,0.4
We move the ToF to 85us, the FLAO Z stage is at 40mm.

Seeing = 1.2"

TN Notes
20141201_074631 G=5.0,0.3,0.1
20141201_074835
20141201_081500 G=1.0,0.3,0.1
We move the FLAO stage to 60mm and we lower the ToF to 83.4us.

TN Notes
20141201_083603 G=1.5,0.5,0.2
20141201_083633
20141201_083844 G=1.5,0.6,0.4
20141201_084032 G=1.5,0.5,0.3
20141201_085238 G=3.0,0.6,0.4
1:00: we optimize the LAN parameters when the FLAO Z stage is at 40mm:

LAN Asti ToF
-185k 1100 83.8
We point again the double star with FLAO bin2. We close the loop on FLAO 150modes to have the proper focus on LUCI. The best focus is at 55mm...

PBL with LAN mirrors their reflecting surface becomes hotter than the backside with the laser propagation...

IMG_20141201_015922180.jpg

The working position for the FLAO DX with IRTC is ~52-54mm.

GLAO on [BDS2003]69

LAN Asti ToF
-160k 900 83.8

We move the ToF to 83us to minimize the LGS slopes values.

We point a single star with FLAO. Z stage is at 50mm, ToF is 84us.

4:00: we have implemented an automatic procedure to offload the LGS focus term on the ToF. Actual value is 83.736us.

TN Notes
20141201_125217 G=8,0.4,0.3, TS off
20141201_125444 ^, TS on
20141201_125519
20141201_125640 ^, seeing=*1.0"*
20141201_125810
20141201_130404 ^, after restart, TS off
20141201_130453 ^, TS on
20141201_130816
20141201_130850
20141201_134138 ^, opening time = 2.0us
20141201_134331
20141201_135058 G=3.0,0.6,0.4
20141201_135618 G=8.0,0.6,0.4
20141201_140955 G=6.0,0.6,0.4

6:00: we lost elevation, at 6:15 we are back

TN Notes
20141201_141804 stuck again
20141201_141941 G=5.0,0.4,0.3
20141201_142144
20141201_142327 G=6.0,0.6,0.4
20141201_142435 G=6.0,0.5,0.4
20141201_143101 G=6.0,0.5,0.4

-- MarcoBonaglia - 18 Dec 2014
Topic revision: r2 - 23 Dec 2014, LorenzoBusoni
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback